# TriSAM: Tri-Plane SAM for zero-shot cortical blood vessel segmentation in VEM images

### 25 Jan 2024 v1

@@ Abstract
In this paper, we address a significant gap in the field of neuroimaging by introducing the largest-to-date public benchmark, BvEM, designed specifically for cortical blood vessel segmentation in Volume Electron Microscopy (VEM) images. The intricate relationship between cerebral blood vessels and neural function underscores the vital role of vascular analysis in understanding brain health. While imaging techniques at macro and mesoscales have garnered substantial attention and resources, the microscale VEM imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure. As researchers delve deeper into the microscale intricacies of cerebral vasculature, our BvEM benchmark represents a critical step toward unraveling the mysteries of neurovascular coupling and its impact on brain function and pathology. The BvEM dataset is based on VEM image volumes from three mammal species: adult mouse, macaque, and human. We standardized the resolution, addressed imaging variations, and meticulously annotated blood vessels through semi-automatic, manual, and quality control processes, ensuring high-quality 3D segmentation. Furthermore, we developed a zero-shot cortical blood vessel segmentation method named TriSAM, which leverages the powerful segmentation model SAM for 3D segmentation. To lift SAM from 2D segmentation to 3D volume segmentation, TriSAM employs a multi-seed tracking framework, leveraging the reliability of certain image planes for tracking while using others to identify potential turning points. This approach, consisting of Tri-Plane selection, SAM-based tracking, and recursive redirection, effectively achieves long-term 3D blood vessel segmentation without model training or fine-tuning. Experimental results show that TriSAM achieved superior performances on the BvEM benchmark across three species.
1 Introduction
With around 2% of body weight, our brain receives around 20% of blood supply. With inadequate blood supply, such as stroke, an average patient loses $\sim$1.9 million neurons (Saver, 2006). Alterations of blood vessel structures are observed in many brain diseases, such as Alzheimer’s and vascular dementia (Kalaria, 2010). Thus, blood vessels in the brain have been extensively investigated by a variety of imaging methods with different resolutions and structural details (Figure 1a). Compared to the macro-level imaging (e.g., CT and MRI (Dyer et al., 2017; McDonald & Choyke, 2003)) and mesoscale-level imaging (e.g.light microscopy Lochhead et al. (2023)), Volume Electron Microscopy (VEM) (Peddie & Collinson, 2014; Peddie et al., 2022) can reveal the detailed ultrastructure including endothelial cells, glial cells, and neurons (Figure 1b).
Traditionally, the imaging methods at the macro and mesoscale are widely used and have produced a large amount of data, and a variety of image segmentation algorithms, public datasets, and evaluation methods have been developed  (Goni et al., 2022; Moccia et al., 2018). Recently, owing to the rapid improvement of imaging technology, the sample size of VEM is significantly increased covering all the layers of the cerebral cortex of mouse (The MICrONS Consortium et al., 2023)and human brain (Shapson-Coe et al., 2021), as well as the whole brain of fly (Dorkenwald et al., 2023). However, to robustly analyze blood vessels in VEM images, existing segmentation methods suffer from two major challenges: the diversity of the image appearance due to variations in the involved imaging pipeline including sample preparation, and the complexity of blood vessel morphology.
@@ 1 Introduction
Figure 1: Comparison of imaging modalities for blood vessel analysis. (a) Both microtomography ($\mu$CT) Dyer et al. (2017) and light microscopy (LM) Lochhead et al. (2023) can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details. (b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all the cell types and their details are visible for further investigations.
@@
To accelerate the method development, we first curate the BvEM dataset, the largest-to-date public benchmark dataset for cortical blood vessel segmentation in VEM images. The VEM image volumes of the BvEM dataset are from recent publications, which are the largest for each of the three mammal species: mouse, macaque, and human acquired at different VEM facilities. We downsampled the volumes to a consistent resolution and performed extensive blood vessel annotation, including manual proofreading, semi-automatic segmentation error correction, and quality control, involving multiple rounds of scrutiny by neuroscience experts to ensure accuracy and completeness.
@@
We then propose a zero-shot 3D segmentation method named TriSAM based on the Segment Anything Model (SAM) (Kirillov et al., 2023) which is able to segment objects in the image given points or a bounding box as the input prompt. With a multi-seed tracking framework, our proposed TriSAM consists of three components: Tri-Plane selection, SAM-based tracking, and recursive redirection. Given the fact that tracking along blood vessel flow direction is easier, our Tri-Plane selection method chooses the best plane for tacking. To further exploit the 3D blood vessel structure and ensure long-term tracking, we propose to change the tracking direction at the potential turning points with a recursive redirection strategy. The proposed method achieves state-of-the-art performance compared to the baseline methods on the proposed BvEM benchmark across all three species.
2 BvEM Dataset
Dataset Design. To foster new methods tackling the challenges, we built our BvEM dataset on top of the largest publicly available VEM image volumes for samples from each of three mammal species: visual cortex from an adult mouse The MICrONS Consortium et al. (2023), superior temporal gyrus from an adult macaque Loomba et al. (2022), and temporal lobe from an adult human Shapson-Coe et al. (2021). Each dataset was acquired with different protocols at different VEM facilities and we refer readers to the respective papers for more details.
Image Volume Standardization. We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and a manageable dataset size for computation and dissemination for AI researchers. Then, we trimmed off the rows and columns across all slices near the image boundary of the mouse and human data, where the blood vessels are hard to annotate due to the missing image content. We directly used the macaque dataset which was well-trimmed by its authors. As shown in Figure 2, the imaging quality and the appearance of blood vessels vary drastically across these three datasets, showcasing the cutting-edge large-scale VEM pipelines in the field.
@@ 2 BvEM Dataset
Figure 2: BvEM Dataset. We compiled the largest publicly available VEM dataset for each of the (a) mouse The MICrONS Consortium et al. (2023), macaque Loomba et al. (2022), and human Shapson-Coe et al. (2021) samples acquired by three different VEM labs. For each dataset, we downsampled the image volume to a similar resolution and trimmed off boundary regions. Then, we corrected the existing ground truth or instance segmentation results for each blood vessel instance (displayed in different colors).
@@
Table 1: Dataset Statistics. Despite the difference in the scale and the geometry of the image volume, the largest blood vessel instance (rendered in red) is significantly bigger than the rest combined.
Sample Microscope Resolution (nm) Volume size (voxel) Length: max/sum (mm)
Mouse TEM 320$\times$ 256$\times$256 2495$\times$3571$\times$2495 1.6/1.7
Macaque SBEM 240$\times$176$\times$176 450$\times$1271$\times$995 713.3/714.5
Human MultiSEM 264$\times$256$\times$256 661$\times$7752$\times$13500 107.2/126.7
Dataset Statistics. We summarize the dataset details in Table 1, where the dimension is in the “zyx" order. The BvEM-Macaque data has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human image data are around 80 and 121 times bigger than it. From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines with the Kimimaro software (Silversmith et al., 2021) and computed the length for each blood vessel instance. Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each dataset. Note that the BvEM-Human dataset is a thin slab, where many blood vessels are disconnected due to the limited spread.
3 Method
Conventional blood vessel segmentation heavily depends on a substantial volume of manually annotated data, a resource that is notably scarce in the existing literature. Therefore, we propose a zero-shot 3D blood vessel segmentation method based on SAM (Kirillov et al., 2023) which achieves superior performance without model training or fine-tuning.
@@ 3 Method
3.1 Problem formulation: 3D Segmentation as multi-seed tracking
To apply SAM to 3D VEM images, a naive method is to combine SAM with a tracking algorithm. However, blood vessels are sensitive to split error requiring robust long-term tracking which is a remaining challenge in object tracking literature. Another possible solution is to use a 3D adaptor with 2D SAM features which requires a large amount of labeled data for training. However, labeled training data is limited since the annotation for VEM images needs the involvement of experts. To this end, we formulate 3D segmentation as a multi-seed tracking framework shown in Figure 3a that doesn’t require any annotated data or robust long-term tracking algorithms. To achieve long-term tracking with multi-seed tracking, we leverage the fact that one of the planes is reliable for tracking while other planes indicate potential turning points. In particular, different from videos, the VEM data can be interpolated from different planes (e.g. yx/zy/zx planes). Tracking along some planes is extremely difficult since the shape and scale change dramatically. Therefore, it is essential to select which plane should be used for tracking. To this end, we propose a Tri-Plane Selection method to select the best plane for tracking based on the potential blood flow direction. To achieve long-term tracking of blood vessels, a Recursive Redirection method is proposed to keep tracking at potential turning points. Based on these two components, a simple SAM-based Tracking is effective in achieving long-term 3D blood vessel segmentation.
3.2 TriSAM 3D Segmentation
The proposed TriSAM 3D segmentation method has three components: Tri-Plane selection, SAM-based tracking, and recursive redirection. The initial seeds are first generated automatically or manually. Then, TriSAM 3D segmentation is performed on each seed and the segmentation results are combined as the final segmentation. The pipeline of the proposed method is shown in Figure 4.
Initial Seed Generation. Initial seeds can be effectively generated with global color thresholding since the pixels of blood vessels are brighter than the background. Therefore, pixels darker than a threshold are considered as background while others are seed candidates. To improve efficiency, we only keep the center of each connected component as the final seeds. In practice, the threshold is set to $\eta I_{m}$ where $I_{m}$ is the maximum intensity of the volume.
@@ 3.2 TriSAM 3D Segmentation
Tri-Plane Selection. Rather than videos, 3D VEM images can be interpreted in different planes. In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically. As shown in Figure 4, tracking along the bottom plane is hard because the shape is complex and changes quickly. This phenomenon can be further confirmed by the experiment shown in Figure 7 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis. Tracking along the plane according to the blood flow direction is ideal and this can be achieved based on choosing a plane with smaller segmentation as shown in Figure 4. Furthermore, we take into account the probability associated with the SAM segmentation result, as it reflects the confidence level in identifying blood vessels. Overall, the plane with the smallest segments and a probability higher than the threshold $\tau$ is selected as the best plane.
SAM-based Tracking. By tracking along the optimally selected plane, changes in shape and size of the blood vessel are more stable compared to other planes. Therefore, we use the enlarged bounding box and the center of the segment as the prompts for the next slice. In particular, we first pad the bounding box of the segment as the bounding box prompt and compute the center of the segment as the point prompt. These prompts are used for segmenting blood vessels in the next slice. The tracking will stop when the probability of the segment is less than a threshold $\tau$.
Figure 4: Left: SAM-based tracking. Right: our proposed TriSAM. (a) Tri-Plane selection is first proposed to select the best plane for tracking. (b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking direction. (c) Recursive redirection exploits potential turning points for long-term tracking.

### 09 Apr 2024 v2

@@ Abstract
While imaging techniques at macro and mesoscales have garnered substantial attention and resources, microscale VEM imaging, capable of revealing intricate vascular details, has lacked the necessary benchmarking infrastructure. In this paper, we address a significant gap in the field of neuroimaging by introducing the largest-to-date public benchmark, BvEM, designed specifically for cortical blood vessel segmentation in volume electron microscopy (VEM) images. Our BvEM benchmark is based on VEM image volumes from three mammal species: adult mouse, macaque, and human. We standardized the resolution, addressed imaging variations, and meticulously annotated blood vessels through semi-automatic, manual, and quality control processes, ensuring high-quality 3D segmentation. Furthermore, we developed a zero-shot cortical blood vessel segmentation method named TriSAM, which leverages the powerful segmentation model SAM for 3D segmentation. To extend SAM from 2D to 3D volume segmentation, TriSAM employs a multi-seed tracking framework, leveraging the reliability of certain image planes for tracking while using others to identify potential turning points. This approach effectively achieves long-term 3D blood vessel segmentation without model training or fine-tuning. Experimental results show that TriSAM achieved superior performances on the BvEM benchmark across three species.
Index Terms:
Blood vessel, Segment anything, Zero-shot segmentation
@@
Figure 1: Comparison of imaging modalities for blood vessel analysis. (a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details. (b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all the cell types and their cellular details.
I Introduction
With around 2% of body weight, our brain receives around 20% of blood supply. Most of the energy and nutrients are consumed by the neurons, and neuron function is sensitive to the blood supply [3, 4]. The blood supply can even be adjusted following the consumption of brain regions, called neurovascular coupling [5]. Alterations of blood vessel structures are observed in many brain diseases, e.g., Alzheimer’s and vascular dementia [6]. Thus, blood vessels in the brain have been extensively investigated with various imaging modalities at different resolutions (Figure 1a). Compared to the macro-level imaging (e.g., CT [1] and MRI [7]) and mesoscale-level imaging (e.g., light microscopy [2]), volume electron microscopy (VEM) [8] can further reveal the detailed ultrastructure including all vascular cells (Figure 1b) for in-depth analysis. However, no large-scale annotated VEM dataset exists to develop and evaluate automated 3D blood vessel segmentation methods.
Traditionally, the imaging methods at the macro and mesoscale are widely used and have produced a large amount of data, and a variety of image segmentation algorithms, public datasets, and evaluation methods have been developed  [9, 10]. At the microscale level, the sample size of VEM is normally limited and most image analyses focus on neuron reconstruction, and blood vessels are largely ignored. Recently, owing to the rapid improvement of imaging technology, the sample size of VEM is significantly increased covering all the layers of the cerebral cortex of mouse [11]and human brain [12], as well as the whole brain of fly [13]. Moreover, imaging the whole mouse brain using VEM technology is under planning [14].
@@ I Introduction
Thus, we first curate the BvEM dataset, the largest-to-date public benchmark dataset for cortical blood vessel segmentation in VEM images to foster segmentation method development. The raw image volumes are from recent publications, which are the largest for each of the three mammal species: mouse, macaque, and human acquired at different VEM facilities. We downsampled the volumes to a consistent resolution and performed extensive blood vessel annotation, including manual proofreading, semi-automatic segmentation error correction, and quality control, involving multiple rounds of scrutiny by neuroscience experts to ensure accuracy and completeness. However, on the BvEM dataset, existing 3D blood vessel segmentation methods suffer from two major challenges: the diversity of the image appearance due to variations in the imaging pipeline and the complexity of the 3D blood vessel morphology. Conventional blood vessel segmentation heavily depends on a substantial volume of manually annotated data, a resource that is notably scarce in the existing literature.
@@
To address these challenges, we propose a zero-shot 3D segmentation method, TriSAM, leveraging the recent 2D segmentation foundation model, Segment Anything Model (SAM) [15], to handle the appearance diversity. To re-purpose SAM for 3D segmentation, we developed a multi-seed segment tracking framework based upon the video object segmentation method with SAM [16]. As it is easier to track 2D blood vessel segments along the blood flow direction, TriSAM selects the best 2D plane for SAM-based tracking and further introduces a recursive seed sampling strategy that performs tri-plane selection at the potential turning points for efficiency. Experimental results demonstrate that the proposed TriSAM method significantly outperforms the prior state-of-the-art methods on the proposed BvEM benchmark across all three species.
II Related Work
Blood Vessel Segmentation. Most existing VEM image segmentation algorithms were developed for neurons [17] and synapses [18, 19, 20]. The traditional VEM dataset size is too small to study blood vessel architecture. Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex [12, 11]. This marks a significant milestone in connectomics research. The blood vessels were segmented manually in the human cortex [12] and automatically using fully convolutional neural network (FCN) [21] or 3D U-Net [11, 22, 23]. Due to the burden of annotation, efforts have been made to decrease the need for annotations [24]. For images acquired using other related imaging methods, such as light microscopy and MRI, many methods were developed and covered by some reviews [9, 10]. Briefly, the blood vessels could be enhanced based on the nature of tubeness using filters like diffusion and ridge filtering. Yet, the segmentation of cortical blood vessels in VEM images is seldom explored due to the absence of publicly available benchmarks.
Evaluation Metrics.. Quantitative evaluation is normally conducted by comparing with manual annotations that are used as the ground truth. The evaluation metrics could be classified into two categories: pixel-based and skeleton-based [10]. Pixel-based metrics measure the difference between segmentation and ground truth at the pixel or voxel level, such as Accuracy and Dice Similarity Coefficient [25]. Skeleton-based metrics measure the difference of skeletons or centerlines, such as Overlap and Overlap Until First Error [26]. Skeleton-based metric requires the blood vessel skeleton that might not always be available. It also requires the computation of point-to-point correspondence between the ground truth and computed skeletons. All existing evaluation metrics are developed using datasets acquired using other imaging methods and we did not find any specific metric for VEM datasets. Most of them could be utilized in VEM images directly. Here we adopt the metrics Precision, Recall, and Accuracy as we measure instance-level pixel prediction.
Segment Anything-Based Models. As a foundation model for image segmentation, the recently proposed Segment Anything Model [15] has garnered significant attention and has been extended to a variety of domains [27, 28, 29, 30] including object tracking [27, 16], image inpainting [28], image mattting [31], super-resolution [32], 3D point cloud [33], and image editting [34]. Despite SAM’s remarkable generalization capabilities, it still encounters some challenges in practical applications, One of these challenges is the huge computation costs due to the heavyweight image encoder. FastSAM [35] adopted a conventional CNN detector with an instance segmentation branch for the segment anything task with real-time speed. MobileSAM [36] proposed decoupled distillation to obtain a small image encoder, which achieved approximately five times faster speed compared to FastSAM while also being seven times smaller in size. Therefore the MobileSAM is employed in our proposed method. One important challenge lies in the unsatisfactory performance of SAM when confronted with special domains, such as medical [37] or biological [38] images, particularly in the context of 3D data. Although SAM demonstrates impressive generalization capabilities, it faces challenges in specialized domains like medical [37, 39] or biological images [38], especially in handling 3D data. [30] assessed the SAM model’s zero-shot segmentation performance in the context of digital pathology and showed scenarios where SAM encounters difficulties. [40] extensively evaluates the SAM for medical image segmentation across 19 diverse datasets, highlighting SAM’s performance variability. To address the domain gap between natural and medical images, SAM-Adapter [37], SAM-Med2D [41], and Medical SAM Adapter [42] introduced Adapter modules and trained the Adapter with medical images. They attained good performance on various medical image segmentation tasks. MedSAM [29] adapted SAM with more than one million medical image-mask pairs and attained accurate segmentation results. MicroSAM [38] also presented a segment anything model for microscopy by fine-tuning SAM with microscopy data. Unlike these approaches that require model fine-tuning for adaptation, our method lifts the blood vessel segmentation capabilities of SAM from 2D images to 3D volumes without any model fine-tuning.
@@ II Related Work
Figure 2: The proposed BvEM dataset. We proofread the blood vessel instance segmentation (displayed in different colors) in the three largest publicly available VEM volumes: (a) mouse [11], macaque [43], and human [12] samples acquired at different VEM labs.
III BvEM Dataset
III-A Sample Information
We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [^1] BvEM-Mouse can be visualized in: https://tinyurl.com/yc2s38x9 [11], superior temporal gyrus from an adult macaque [43], and temporal lobe from an adult human [12]. Each volume was acquired with different protocols at different facilities and we refer to respective papers for more details.
TABLE I: Dataset information. Despite the difference in the scale and the geometry of the image volume, the largest blood vessel instance (rendered in red) is significantly bigger than the rest combined. The last column shows the max/total length of the blood vessel instances.
Sample Resolution (nm) Size (voxel) Length: (mm)
Mouse (TEM) 320$\times$ 256$\times$256 2495$\times$3571$\times$2495 1.6/1.7
Macaque (SBEM) 240$\times$176$\times$176 450$\times$1271$\times$995 713.3/714.5
Human (MultiSEM) 264$\times$256$\times$256 661$\times$7752$\times$13500 107.2/126.7
III-B Image Volume Standardization
We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and a manageable dataset size for AI researchers. Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content. As shown in Fig. 2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
III-C Blood Vessel Instance Segmentation Annotation
@@
Figure 4: TriSAM framework. (a) Tri-plane selection is first proposed to select the best plane for tracking. (b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking direction. (c) Recursive seed sampling exploits potential turning points for long-term tracking.
III-D Dataset Statistics
We summarize the dataset details in Tab. I. The BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively. From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [45] and computed the length for each blood vessel instance. Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume. The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3 where the dimension is in the “zyx” order.
IV Method
IV-A Overview
We formulate the 3D blood vessel segmentation task as a multi-seed video object segmentation task: generating one 3D segment per seed through tracking 2D SAM results along one fixed direction and the final output is the connected component of the combined segmentation results (Fig. 5a). However, unlike videos, the 3D blood vessel segment can be tracked from different axes, whereas 2D SAM segmentation results can be of drastically different qualities. As shown in Fig. 5b, the seed generation can have many errors leading to incomplete segmentation results. To achieve long-term tracking, the framework needs a recursive seed sampling module to propose potential turning points as new seeds based on segmentation results. As shown in Fig. 5c, the ground truth segment on some 2D planes can have such an unnatural shape that SAM fails to segment it, not to mention its drastic shape change along the axis. Ideally, one axis is along the blood flow direction where the 2D tracking plane is an ellipse shape, e.g., zx plane in the example. Thus, it is essential to have a tri-plane selection module in the framework.
The proposed TriSAM framework has the following three modules: tri-plane selection, SAM-based tracking, and recursive seed sampling (Fig. 4).
@@ IV-A Overview
IV-B Tri-Plane Selection
@@
Rather than videos, 3D VEM images can be interpreted in different planes. In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically. This phenomenon can be further confirmed by the experiment shown in Table III in which tracking along the $y$ axis is more effective than tracking along the $x$ axis. Given the seed position as a prompt, we run SAM on each of the tri-planes of the 3D image volume centered at the seed. Note that the plane that is most tangent to the blood flow direction has the smallest segment size, capturing the blood vessel’s cross-section (Fig. 4a, xy-plane). In addition, SAM outputs the probability for the segment result, reflecting the confidence of the naturalness of its shape. Thus, to combine the above cues for tri-plane selection, we pick the 2D plane with the smallest segment size and a probability higher than the threshold $\tau$. TriSAM terminates when no plane is selected.
IV-C SAM-based Tracking
Given the segment on the selected 2D plane from above, we aim to track it with SAM along the orthogonal axis. If we use the center of the current segment as the point prompt for the next slice, SAM may output segments with inconsistent sizes along the axis due to the lack of size information in the point prompt. If we use the bounding box of the current segment as the box prompt, SAM may not track in the long term as the output enclosed in the bounding box is getting smaller and smaller. Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results. TriSAM terminates when SAM returns no prediction for the next image slice.
IV-D Recursive Seed Sampling
The above modules work well when the blood vessel travels in one direction. In reality, blood vessels not only turn in different directions but also bifurcate, leading to drastic shape and scale changes as shown in Fig. 11. Thus, we need a module to propose potential turning points to recursively run TriSAM, linking short-term segment tracks into long-term segment tracks. SAM-based tracking is simple and fast for short-term tracking. But it may fail when shape and scale change dramatically as shown in Fig. 11. Therefore, it relies on the large number of seeds which is hard to generate as shown in Fig. 5(b).
To enable long-term tracking, we propose a recursive seed sampling method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels. Given the 3D segment from the short-term track after TriSAM terminates, we first use the Farthest Point Sampling (FPS) to find potential endpoints. Then, we add these points to the global seed tree whose root note is the input seed. To avoid repeating the track along the same direction as the previous short-term track, these newly sampled seeds store the previous tracking direction. For example, if the previous track is along the z-axis, TriSAM on the new seed will not compute results along the z-axis to encourage exploration of new tracking directions. During each iteration, we take an unvisited node in the seed tree, mark it as visited, and run TriSAM along a different tracking direction. With such a seed tree data structure, we run TriSAM recursively until all potential points are visited. For each endpoint, we re-run the above two TriSAM modules to track segments along directions that are different from the current tracking axis.
Mathematically, we denote the 3D coordinate of a turning point as ($x_{t}$,$y_{t}$,$z_{t}$). At this point, we perform the SAM operation on three planes separately. We use $M_{z}(x,y,z_{t})$, $M_{y}(x,y_{t},z)$, and $M_{x}(x_{t},y,z)$ to represent the segmented mask region along xy plane, xz plane, and yz plane, respectively. Assuming that the chosen best plane is xz plane which means we will track along the y-axis, we will obtain new turning points based on the segmented masks on the non-best planes $M_{z}(x,y,z_{t})$ and $M_{x}(x_{t},y,z)$. Since the boundary region of the mask corresponds to where blood vessels disappear, it indicates locations where there is a change in the direction of blood vessel flow. So, the boundaries in the y-direction of both $M_{z}(x,y,z_{t})$ and $M_{x}(x_{t},y,z)$ indicate locations where blood vessels change direction along the y-axis. We further define the maximum and minimum values of the y-coordinate within the $M_{z}(x,y,z_{t})$ region as $y_{max}^{z_{t}}$ and $y_{min}^{z_{t}}$, respectively. Similarly, we can define $y_{max}^{x_{t}}$ and $y_{min}^{x_{t}}$. Then the new turning points can be defined as $\{(x,y,z)|(x,y,z)\in M_{z}(x,y_{max}^{z_{t}},z_{t})\cup M_{z}(x,y_{min}^{z_{t}},z_{t})\cup M_{x}(x_{t},y_{max}^{x_{t}},z)\cup M_{x}(x_{t},y_{min}^{x_{t}},z)\}$. The turning point will be dropped if it hits the volume boundary or the probability of its segment is less than the threshold $\tau$. After updating the turning point tree, we then sample a new leaf node for the next iteration until all leaf nodes are sampled.